# How to set AdamW’s weight decay as you scale model and dataset size

### v1

@@ 5 Related work
Our EMA view of AdamW provides an explanation for several striking observations made in the literature.
@@
First, the original AdamW paper (Loshchilov & Hutter, 2018) proposed a parameterization in terms of the learning rate, $\eta$ and $\gamma=\eta\lambda$ describing the weight decay. They made an empirical observation that $\eta$ and $\gamma$ were more decoupled than $\eta$ and $\lambda$ (in the sense that the optimal $\eta$ depended strongly on $\lambda$ and less strongly on $\gamma$). We are able to provide a theoretical understanding of their observation, by noting that $\gamma$ is related to our timescale, $\gamma=\nicefrac{{1}}{{\tau_{\text{iter}}}}$. We are therefore able to identify two separate processes one associated with $\eta$ and another associated with $\gamma=\nicefrac{{1}}{{\tau_{\text{iter}}}}$. First, $\eta$ describes the size of the initial updates relative to the random initialization, which is relevant close to initialization (see Appendix A for details). In contrast, $\gamma=\nicefrac{{1}}{{\tau_{\text{iter}}}}$ describes the EMA timescale, which is relevant as the model approaches convergence.
@@
Second, this viewpoint suggests that the effects of the learning rate, $\eta$, on the magnitude of the updates relative to the random initialization should become less relevant than the EMA timescale as training proceeds (Appendix A). While this seems strange (of course the learning rate matters alot), there are some suggestions in the literature that indeed the EMA timescale may be more important in some settings. Specifically, Wortsman et al. (2024) ran proxies for large-scale LLM pretraining. They showed that if $\gamma=\eta\lambda=\nicefrac{{1}}{{\tau_{\text{iter}}}}$ is fixed, the final validation loss is relatively insensitive to the learning rate, $\eta$. In contrast, if $\lambda$ is fixed, the validation loss shows much more pronounced sensitivity to the learning rate, $\eta$. This makes sense if the key hyperparameter is $\tau_{\text{iter}}$, as modifying the learning rate, $\eta$, with $\lambda$ fixed implies $\tau_{\text{iter}}=1/(\eta\lambda)$ also changes.
@@
Finally, µP itself is related work (Yang et al., 2022), and is discussed extensively in the Background section. Additionally, the $\gamma=\eta\lambda$ parameterization is often used in SGD, where it is identified as the effective learning rate (Van Laarhoven, 2017; Zhang et al., 2019; Li & Arora, 2019; Li et al., 2020; Wan et al., 2021; Li et al., 2022), though this literature is not clear on whether those results transfer to Adam or AdamW. Finally, recent work (Busbridge et al., 2023) studies the scenarios where an EMA is used for averaging model weights, such as in self-supervised learning, and proposes hyperparameter scaling rules to achieve the same performance under different batch sizes. Our work differs in that we are interpreting AdamW as EMA over past updates rather than using explicitly using EMA for averaging weights.
6 Limitations
There are other update rules such as Lion (Chen et al., 2023) and Sophia (Liu et al., 2024) that use decoupled weight decay. We expect that our recommendations will transfer to those settings, though we have not investigated this explicitly as AdamW is by far and away the most popular method at present.
@@ 7 Conclusions
We showed that AdamW’s weight updates can be understood as an EMA. The critical EMA hyperparameter is the EMA timescale, $\tau_{\text{epoch}}$, and we can use $\tau_{\text{epoch}}$ (in combination with the learning rate, $\eta$) to set the weight decay. We gave guidelines for setting $\tau_{\text{epoch}}$, and argued that $\tau_{\text{epoch}}$ should remain fixed as we change model and dataset size. This implies that the optimal weight decay should decrease with dataset size and increase with model size, and we validated these predictions experimentally.
@@ Appendix A Confirming the connection between AdamW and weight decay
The key insight from the EMA view on AdamW is that an EMA (Eq. 5), has only two choices, the EMA timescale, $\tau_{\text{iter}}$, and the initialization, $\text{ema}_{0}$. Here, we prove that there are similarly only two choices when using AdamW in a scale-invariant network (see Eq. 17). In particular, we show that for a carefully chosen initialization, the full trajectory of weights optimized by AdamW is controlled by $\tau_{\text{iter}}$ and the size of the first step, relative to the initial parameter scale $\rho=\eta/\sigma$.
Theorem 1.
Consider two AdamW optimizers with different learning rates and weight decays, and initialization scales, ($\eta,\lambda,\sigma$ vs. $\eta^{\prime},\lambda^{\prime},\sigma^{\prime}$). Take $w_{t}$ to be the parameters learned by the first optimizer after the $t$th optimization step, and $w_{t}^{\prime}$ to be the parameters learned by the second optimizer. These optimizers are initialized by
$w_{0}$ $=\sigma{\xi}$ (16a)
$w_{0}^{\prime}$ $=\sigma^{\prime}{\xi},$ (16b)
where $\boldsymbol{\xi}$ is random noise (e.g. IID Gaussian).
Consider a scale-invariant network, in the sense that multiplying the weights, $w$, by an arbitrary positive constant, $c>0$ gives the same output for all inputs, $x$,
$\operatorname{net}(x;w)$ $=\operatorname{net}(x;\tfrac{1}{c}w).$ (17)
We use the same EMA timescale,
$\nicefrac{{1}}{{\tau_{\text{iter}}}}=\eta\lambda=\eta^{\prime}\lambda^{\prime},$ (18)
and the same ratio of the first step, relative to the initial parameter scale,
$\rho=\tfrac{\eta}{\sigma}=\tfrac{\eta^{\prime}}{\sigma^{\prime}}.$ (19)
Now, the entire trajectory of weights for each optimizer is the same (up to a scalar multiplier,)
$w_{t}^{\prime}$ $=\tfrac{1}{c}w_{t},\forall t$ (20)
where
$c$ $=\tfrac{\sigma}{\sigma^{\prime}}=\tfrac{\eta}{\eta^{\prime}}=\tfrac{\lambda^{\prime}}{\lambda}.$ (21)
Thus, the network outputs are the same at all points along the trajectory,
$\operatorname{net}(x;w_{t})$ $=\operatorname{net}(x;w_{t}^{\prime}).$ (22)
@@ A.1 Proof of Theorem 1
Here, we consider two AdamW training trajectories with hyperparameters $\eta,\lambda,\epsilon$ and $\eta^{\prime},\lambda^{\prime},\epsilon^{\prime}$. We prove that for a specific way of initializing the optimizer state, and if the optimizer hyperparameters are related by,
$\eta^{\prime}$ $=\tfrac{1}{c}\eta$ (23)
@@
then the full trajectories are the same. Notice that $\nicefrac{{1}}{{\tau_{\text{iter}}}}=\eta\lambda=\eta^{\prime}\lambda^{\prime}$ remains unchanged with this choice of hyperparameters.
@@
$w_{t}$ $=(1-\lambda\eta)w_{t-1}+\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (26f)
(And the analogous equations for the primed optimizer). The optimizer state is the variables that persist across timesteps, namely $m$, $v$, and $w$. Specifically, the first optimizer has state $m_{t}$, $v_{t}$, and $w_{t}$ while the second optimizer has state $m^{\prime}_{t}$, $v^{\prime}_{t}$, and $w^{\prime}_{t}$. Our goal is to prove that for a scale-invariant network, as $\eta$ or $\lambda$ changes but $\nicefrac{{1}}{{\tau_{\text{iter}}}}=\eta\lambda$ is fixed, the states for the different optimizers are related by scale parameters,
@@
where $c=\lambda^{\prime}/\lambda=\eta/\eta^{\prime}$. The argument proceeds inductively, by assuming that the scaling relations in Eq. (27) hold at timestep t1t-1, then proving that as a consequence they hold at timestep $t$.
@@
$w_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w_{t-1}+\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (34a)
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w^{\prime}_{t-1}+\eta^{\prime}\frac{\hat{m}^{\prime}_{t}}{\sqrt{\hat{v}^{\prime}_{t}}+\epsilon^{\prime}}$ (34b)
Substituting $\eta=\nicefrac{{1}}{{\tau_{\text{iter}}}}/\lambda$ and $\eta^{\prime}=\nicefrac{{1}}{{\tau_{\text{iter}}}}/\lambda^{\prime}$,
$w_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (35a)
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w^{\prime}_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{\lambda^{\prime}}\frac{\hat{m}^{\prime}_{t}}{\sqrt{\hat{v}^{\prime}_{t}}+\epsilon^{\prime}}$ (35b)
@@
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w^{\prime}_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{c\lambda}\frac{c\hat{m}^{\prime}_{t}}{\sqrt{c^{2}\hat{v}^{\prime}_{t}}+c\epsilon}$ (36)
@@
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})\tfrac{1}{c}w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{c\lambda}\frac{c\hat{m}_{t}}{\sqrt{c^{2}\hat{v}_{t}}+c\epsilon}$ (37)
@@
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})\tfrac{1}{c}w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{c\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}}$ (38)
@@
$w^{\prime}_{t}$ $=\tfrac{1}{c}\left((1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}}\right)=\tfrac{1}{c}w_{t}.$ (39)
@@
Thus, we have shown that if the scaling relations in Eq. (27) hold at timestep t1t-1, they also hold at timestep $t$. It only remains to complete the inductive proof by showing that the scaling relations hold at initialization. The $m$ and $v$ state variables are usually initialized at zero, $m_{0}=v_{0}=0$, and the scaling relations of course hold at zero. However, networks are usually initialized with a fixed parameter scale which does not depend on $\lambda$. To ensure full scale-invariant training, we therefore need to rescale the parameter initializations, to ensure $w_{0}^{\prime}=\tfrac{1}{c}w_{0}=\tfrac{\eta^{\prime}}{\eta}w_{0}$.
@@ A.3 Empirical confirmation of Theorem 1
[…]
This suggests that standard network setups require considerable modification before training trajectories actually become invariant to $\eta$ for a fixed $\nicefrac{{1}}{{\tau_{\text{iter}}}}$ (or $\tau_{\text{epoch}}$). In particular, recall that Theorem 1 has two major assumptions about the model: 1. It needs to be scale-invariant (Eq. 17); 2. The initialization needs to be $\eta$-aware (Eq. 19). As such, we need to make three key modifications in order for the assumptions to hold:
@@
We additionally considered ViT under a similar setting. Notice that ViT needs more extensive modifications to ensure full scale-invariance; full details are presented in Appendix. A.3.4. The performance vs. iteration and timescale for various versions of the model are presented in Fig. 8 and 9. Mirroring the ResNet results, we again found that, under the suggested modifications, the performance becomes invariant to $\eta$ under fixed $\tau_{\text{epoch}}$.
@@
[…]

### v2

@@ 5 Related work
There have been a number of follow-up works that follow up on the concepts presented in this paper and cite us. First, Blake et al. (2024) confirmed that increasing weight decay as we increase model width fixes learning rate transfer. Second, Anonymous (2024) has built on our EMA interpretation of AdamW to understand and propose better learning rate schedules. Third, Bjorck et al. (2024) looked at the scaling of optimal learning rates with dataset size. They found that the optimal learning rate decreases with dataset size ($N$), with a fixed batch size, and a fixed weight decay. This broadly what you would expect from our findings, that you need to increase the EMA timescale as the dataset size increases, and you can do that by reducing the weight decay, or reducing the learning rate. However, they found that the dependence of learning rate on dataset size was sublinear. We suspect this is because as you modify the learning rate with weight decay fixed, you simultaneously change two things: the EMA timescale (which you probably do want to change) and the behaviour of the optimizer close to initialization, which you probably do not want to change (Appendix A).
@@
Further, our EMA view of AdamW provides an explanation for several striking observations made in the literature.
@@
First, the original AdamW paper (Loshchilov & Hutter, 2018) proposed a parameterization in terms of the learning rate, $\eta$ and $\gamma=\eta\lambda$ describing the weight decay. They, as well as a recent blog post (Schaipp, 2024), made an empirical observation that $\eta$ and $\gamma$ were more decoupled than $\eta$ and $\lambda$ (in the sense that the optimal $\eta$ depended strongly on $\lambda$ and less strongly on $\gamma$). We are able to provide a theoretical understanding of their observation, by noting that $\gamma$ is related to our timescale, $\gamma=\nicefrac{{1}}{{\tau_{\text{iter}}}}$. We are therefore able to identify two separate processes one associated with $\eta$ and another associated with $\gamma=\nicefrac{{1}}{{\tau_{\text{iter}}}}$. First, $\eta$ describes the size of the initial updates relative to the random initialization, which is relevant close to initialization (see Appendix A for details). In contrast, $\gamma=\nicefrac{{1}}{{\tau_{\text{iter}}}}$ describes the EMA timescale, which is relevant as the model moves away from the initialization. Finally, Loshchilov & Hutter (2018), did not use these insights to propose how to scale the weight decay with model and dataset size (our key contribution).
@@
Second, our viewpoint (Sec. 3.2) suggests that the effects of the learning rate, $\eta$, on the magnitude of the updates relative to the random initialization should become less relevant than the EMA timescale as training proceeds (Appendix A). While this seems strange (of course the learning rate matters alot), there are some suggestions in the literature that indeed the EMA timescale may be more important in some settings. Specifically, Wortsman et al. (2024) ran proxies for large-scale LLM pretraining. Fig. 6 in their paper showed that if $\gamma=\eta\lambda=\nicefrac{{1}}{{\tau_{\text{iter}}}}$ is fixed, the final validation loss is relatively insensitive to the learning rate, $\eta$. In contrast, if $\lambda$ is fixed, the validation loss shows much more pronounced sensitivity to the learning rate, $\eta$. This makes sense if the key hyperparameter is $\tau_{\text{iter}}$, as modifying the learning rate, $\eta$, with $\lambda$ fixed implies $\tau_{\text{iter}}=1/(\eta\lambda)$ also changes. Indeed, the $\gamma=\eta\lambda$ parameterization is often used in SGD, where it is identified as the effective learning rate (Zhang et al., 2019b; Li & Arora, 2019; Li et al., 2020; Wan et al., 2021; Li et al., 2022). Recent works extend the effective learning rate perspectives to AdamW by viewing AdamW as sign gradient descent (D’Angelo et al., 2024) or as rotation speed on a fixed radius sphere (Kosson et al., 2023). However, these literautres do not study how optimal $\eta\lambda$ transfers across model and dataset sizes.
@@
Fourth, many papers have considered scaling with batch size (which we hold fixed) (Zhang et al., 2019a; Malladi et al., 2022; Wang & Aitchison, 2024). These papers find that when the batch size is small, the optimal learning rate is also small. Then, as the batch size increases, the optimal learning rate also increases. However, there’s a limit, the critical batch size, at which point increases in the batch size no longer allow you to increase the learning rate. Once you hit the critical batch size, you no longer get any benefit from increasing the batch size, and you instead must increase the number of iterations (McCandlish et al., 2018; Shallue et al., 2019). Note that while very recent work suggests that the optimal batch size does increase with dataset size, they confirm that this increase is sublinear, implying that the number of training iterations will increase (Zhang et al., 2024). As soon as you need to increase the number of iterations, our recommendation to hold $\tau_{\text{epoch}}$ constant comes into play.
Concurrent work (D’Angelo et al., 2024) suggests that weight decay does not serve as regularization in deep learning, but rather controls minibatch noise, which grows with $\eta\lambda$. Our work provides a similar viewpoint, in the sense that, if you are averaging over more data points (higher $\tau_{\text{iter}}=1/\eta\lambda$), then the iteration-to-iteration noise will of course be smaller. But the EMA framework also provides you with guidelines for how to set $\lambda$ or equivalently the timescale as you scale model and dataset size.
Of course, µP itself is related work (Yang et al., 2022), and is discussed extensively in the Background section. Van Laarhoven (2017) argue that in Adam, the weight decay hyperparameter changes the scale of the learned weights, and thereby has an influence on the effective learning rate. This effect pops up in our framework in the $1/\lambda$ factor in $q_{t}$ (Eq. 6), the quantity over which we take the EMA. Additionally, recent work (Busbridge et al., 2023) studies the scenarios where an EMA is used for averaging model weights, such as in self-supervised learning, and proposes hyperparameter scaling rules to achieve the same performance under different batch sizes. Our work differs in that we are interpreting AdamW as EMA over past updates rather than using explicitly using EMA for averaging weights. Finally, hints that the optimal weight decay decreases with dataset size have appeared previously in empirical literature when the authors have conducted thorough hyperparameter sweeps, though they did not propose a quantitative, theoretically motivated scaling rule (Zhai et al., 2019).
@@ 6 Conclusions
We showed that AdamW’s weight updates can be understood as an EMA, showed that the optimal EMA timescale is fixed as we scale model and dataset size, and explored the implications for hyperparameter scaling of the weight decay.
@@ A.1 Proof of Theorem 1
Here, we consider two AdamW training trajectories with hyperparameters $\eta_{t},\lambda,\epsilon$ and $\eta_{t}^{\prime},\lambda^{\prime},\epsilon^{\prime}$. We prove that for a specific way of initializing the optimizer state, and if the optimizer hyperparameters are related by,
$\eta_{t}^{\prime}$ $=\tfrac{1}{c}\eta_{t}$ (17)
@@
then the full trajectories are the same. Notice that $\nicefrac{{1}}{{\tau_{\text{iter};t}}}=\eta_{t}\lambda=\eta_{t}^{\prime}\lambda^{\prime}$ remains unchanged with this choice of hyperparameters.
@@
$w_{t}$ $=(1-\lambda\eta_{t})w_{t-1}+\eta_{t}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (20f)
(And the analogous equations for the primed optimizer). The optimizer state is the variables that persist across timesteps, namely $m$, $v$, and $w$. Specifically, the first optimizer has state $m_{t}$, $v_{t}$, and $w_{t}$ while the second optimizer has state $m^{\prime}_{t}$, $v^{\prime}_{t}$, and $w^{\prime}_{t}$. Our goal is to prove that for a scale-invariant network, as $\lambda$ or the trajectory of $\eta_{t}$ changes while the trajectory of $\nicefrac{{1}}{{\tau_{\text{iter};t}}}=\eta_{t}\lambda$ is fixed, the states for the different optimizers are related by scale parameters,
@@
where $c=\lambda^{\prime}/\lambda=\eta_{t}/\eta_{t}^{\prime}$. The argument proceeds inductively, by assuming that the scaling relations in Eq. (21) hold at timestep $t-1$, then proving that as a consequence they hold at timestep $t$.
@@
$w_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter};t}}})w_{t-1}+\eta_{t}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (28a)
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter};t}}})w^{\prime}_{t-1}+\eta_{t}^{\prime}\frac{\hat{m}^{\prime}_{t}}{\sqrt{\hat{v}^{\prime}_{t}}+\epsilon^{\prime}}$ (28b)
Substituting $\eta_{t}=\nicefrac{{1}}{{\tau_{\text{iter};t}}}/\lambda$ and $\eta^{\prime}=\nicefrac{{1}}{{\tau_{\text{iter};t}}}/\lambda^{\prime}$,
$w_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter};t}}})w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter};t}}}\frac{1}{\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (29a)
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter};t}}})w^{\prime}_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter};t}}}\frac{1}{\lambda^{\prime}}\frac{\hat{m}^{\prime}_{t}}{\sqrt{\hat{v}^{\prime}_{t}}+\epsilon^{\prime}}$ (29b)
@@
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter};t}}})w^{\prime}_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter};t}}}\frac{1}{c\lambda}\frac{c\hat{m}^{\prime}_{t}}{\sqrt{c^{2}\hat{v}^{\prime}_{t}}+c\epsilon}$ (30)
@@
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter};t}}})\tfrac{1}{c}w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter};t}}}\frac{1}{c\lambda}\frac{c\hat{m}_{t}}{\sqrt{c^{2}\hat{v}_{t}}+c\epsilon}$ (31)
@@
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter};t}}})\tfrac{1}{c}w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter};t}}}\frac{1}{c\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}}$ (32)
@@
$w^{\prime}_{t}$ $=\tfrac{1}{c}\left((1-\nicefrac{{1}}{{\tau_{\text{iter};t}}})w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter};t}}}\frac{1}{\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}}\right)=\tfrac{1}{c}w_{t}.$ (33)
@@
Thus, we have shown that if the scaling relations in Eq. (21) hold at timestep $t-1$, they also hold at timestep $t$. It only remains to complete the inductive proof by showing that the scaling relations hold at initialization. The $m$ and $v$ state variables are usually initialized at zero, $m_{0}=v_{0}=0$, and the scaling relations of course hold at zero. However, networks are usually initialized with a fixed parameter scale which does not depend on $\lambda$. To ensure full scale-invariant training, we therefore need to rescale the parameter initializations, to ensure
$w_{0}^{\prime}=\tfrac{1}{c}w_{0}=\tfrac{\eta^{\prime}}{\eta}w_{0}.$ (34)
@@ A.3 Empirical confirmation of Theorem 1
In this section, we empirically confirm Theorem 1 using a ResNet-18 and a ViT (with QK-layernorm) trained on CIFAR-10. We used AdamW to perform the optimization for 200 epochs with a cosine learning rate schedule to decay the initial learning rate by a factor of 10. We used a fixed batch size of 100 and tested initial learning rates $\eta$ in range $\{10^{-6},10^{-5},\ldots,10^{-1}\}$. For each $\eta$, we considered $\lambda$ in range $\{10^{-7}/\eta,10^{-6}/\eta,\ldots,10^{-2}/\eta\}$, giving us $\tau_{\text{epoch}}$ in range $\{2\times 10^{-1},2\times 10^{0},\ldots 2\times 10^{4}\}$ for all $\eta$s.
We begin by looking at the performance of ResNet under different values of $\tau_{\text{epoch}}$ and $\eta$ under the standard configuration. Fig. 7B and Fig. 8B show performance metrics vs. timescale for different learning rates after 200 epochs. We can see from Fig. 7B that the behavior for different learning rates, $\eta$, is very different. In addition, Fig. 8B shows that, while for most $\eta$s, the optimal $\tau_{\text{epoch}}$ in terms of test metrics is the same, the exact performance at the optimal $\tau_{\text{epoch}}$ differs.
This suggests that standard network setups require considerable modification before training trajectories actually become invariant to $\eta$ for a fixed $\nicefrac{{1}}{{\tau_{\text{iter}}}}$ (or $\tau_{\text{epoch}}$). In particular, recall that Theorem 1 has two major assumptions about the model: 1. It needs to be scale-invariant (Eq. 9); 2. The initialization needs to be $\eta$-aware (Eq. 34). Therefore, we need to make three key modifications in order for the assumptions to hold:
@@
We additionally considered ViT under a similar setting. Notice that ViT needs more extensive modifications to ensure full scale-invariance; full details are presented in Appendix. A.3.4. The performance vs. iteration and timescale for various versions of the model are presented in Fig. 9 and 10. Mirroring the ResNet results, we again found that, under the suggested modifications, the performance becomes invariant to $\eta$ under fixed $\tau_{\text{epoch}}$.
@@
Figure 10: Under proposed modifications, ViT shows performance controlled only by the timescale and irrelevant to the learning rate. Similar to Fig. 9 but now we plotted the final performance metrics against $\tau_{\text{epoch}}$ (x-axis) under different $\eta$ (lines). In rows B, and C, the model is not fully scale-invariant and the initialization is not $\eta$-dependent, as such the performance varies between $\eta$s under a fixed $\tau_{\text{epoch}}$. In row D, when all modifications are incorporated, i.e. assumptions in Theorem. 1 are satisfied, the performance becomes only dependent on $\tau_{\text{epoch}}$.